Exercise 26.44 14 of 26 Part A Constants A 12.0 ?F capacitor is connected through a 0.885 ?? resistor to a constant potential difference of 60.0 v. Compute the charge on the capacitor at the following times after the connections are made: 0, 5.0 s, 10.0 s, 20.0 s, and 100.0s Express your answers using two significant figures. Enter your answers numerically separated by commas. Submit ? Part B Compute the charging currents at the same instants Express your answers using two significant figures. Enter your answers numerically separated by commas. io, i5.0, iio, i20, i1oo Submit
Answer #1

Time constant for the circuit, ( 0.885x 106 Ω)(12x10-6 F ) -10.62 s Charge at the capacitor is, (12x10-6)(60)| 1-er 720x10-01Part A: q = 720 x 10-6 | 1-e 10.62 Yo = 720 x 10-6 | 1-e-10.62 720x10-1-e T0.62 2.699 × 10 = 2.7×10-4 C 10 Ao = 720x10-6 | 1-Part b Current at instant t 0.885x106 = 67.79 x10-61 e-- 1 = 67.79 x10ー‘l e 10.62 67.79x1010.62 67.79x10-5 = 68 × 10-6 A 67.7